The firm is a monopsonist in the labor market and a price taker in the output market. Labor demand is l^D=12 (i.e. every worker has a constant MRP_l of 12). Labor supply is (w)=square root of w. The government imposes a minimum wage of w=12. What is the wage rate in this economy? Enter a number only.
Answer #1

Let us calculate the equilibrium wage rate before minimum wage regulation.

Given that

LD=12 (MRPL=12)

Labor supply is given by

L=w0.5 or w=L2

Total Cost=TC=L*w=L2*L=L3

Marginal Cost of labor=MCL=dTC/dL=3L2

Set MCL=MRPL

3L2=12

L=2

w=22=4

Siince, minimum wage is higher than equilibrium wage rate. So, it is binding.

So, Wage rate in this economy=minimum wage rate=12